Here is the problem I am trying to solve:
The sum of the digits of a two-digit number is 6. If the digits are reversed, the new number is 9 less than four times the original number. Find the original number.
Thanks!

Answer by mananth(16949)

let the number be xy. x in the tens place.

x+y=6------------(1)
10y+x=4(10x+y)-9
10y+x=40x+4y-9
6y-39x=-9------------(2)
Add equation (1) & (2)
1.00 x + 1 y = 6 .............1
-39 x + 6 y = -9 .............2
Eliminate y
multiply (1)by -6
Multiply (2) by 1
-6 x -6 y = -36
-39 x + 6 y = -9
Add the two equations
-45 x = -45.000
/ -45
x = 1
plug value of x in (1)
x + y = 6
+ y = 6
y = 6 -1
y = 5
y = 5
The number is 15
m.ananth@hotmail.ca